You’ve observed the following returns on Yasmin Corporation’s stock over the past five years: 20 percent, –12 percent, 17 percent, 20 percent, and 10 percent. Suppose the average inflation rate over this period was 1.7 percent and the average T-bill rate over the period was 4.6 percent.

a. What was the average real return on the company's stock? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

Average real return ______   %

b. What was the average nominal risk premium on the company's stock? (Do not round intermediate calculations. Enter your answer as a percent rounded to 1 decimal place, e.g., 32.1.)

Average nominal risk premium _______ %
Answer #1

a

Average return = (.20-.12+.17+.20+.10) / 5=11.00%

To calculate the average real return, we can use the average return of the asset, and the average inflation in the Fisher equation.

(1 + R) = (1 + r)(1 + h)
= (1+11%)/(1+1.7%) - 1

Average real return=9.14%

b

Average risk premium = Average return − Average risk-free rate

= .110 − .046=.0640

Average risk premium=6.40%